In a series connection solve for the total capacitance.


C1 = 3.5 x 10^-12 F

C2 = 2.41 x 10^-12 F

C3 = 1.13 x 10^-12 F

V = 1700 V


1
Expert's answer
2021-11-01T12:54:29-0400

Solution;

Capacitance in series is given as;

"\frac{1}{C_T}=\frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3}+...."

By substitution;

"\frac{1}{C_T}=\frac{1}{3.5×10^{-12}}+\frac{1}{2.41×10^{-12}}+\frac{1}{1.13×10^{-12}}"

"\frac{1}{C_T}=1.5856×10^{12}"

"C_T=\frac{1}{1.5856×10^{-12}}"

"C_T=6.307\u00d710^{-13}" F